OpenStudy (anonymous):

Calculus II: Starting at s = 0 when t = 0, an object moves along a line so that its velocity at time t is v(t) = 2t - 4 centimeters per second. How long will it take to get to s = 12? To travel a total distance of 12 centimeters?

OpenStudy (anonymous):

I have t=6

OpenStudy (anonymous):

hold on, i'll compute and see if it matches

OpenStudy (anonymous):

v(t) = 2t - 4 s(t) = integral 2t - 4 = 2t^2 - 4t + C 0 = 0^2 -4(0) + C t^2 - 4t = 12 t^2 -4t -12=0 (t-6)(t-2) t= 6 considering time is never negative

OpenStudy (anonymous):

Grrrrrrr Idk HHHHEEEEELLLLLPPPPPP

OpenStudy (anonymous):

that will be (t-6)(t+2)

OpenStudy (anonymous):

yes i get 6 also

OpenStudy (anonymous):

second thing, u have made the displacement 0, but actually distance is needed

OpenStudy (anonymous):

he is correct

OpenStudy (anonymous):

sorry, u have made displacement 12 :)

OpenStudy (anonymous):

yes since the velocity starts out negative, the object is going backwards at first, then goes forwards. do you want the total displacement to be 12 or the position from the starting point to be 12?

OpenStudy (anonymous):

Arnab09 would that mean I would have to integrate as such